Reports 1-1 of 1 Clear search Modify search
AdV-TCS (CO2 laser projector)
gherardini, melo, nardecchia - 18:24 Tuesday 26 August 2025 (67569) Print this report
Damaged mirror replacement on the NI bench / Part 2

This afternoon, Suzanne and Fabio adjusted the M3 mirror to restore the pre-mirror replacement alignment.

The comparison of the DAS before and after the mirror replacement is shown in Fig. 1.

Work on the bench was completed at 15:55 UTC.
The correct powers were recovered, and all flip mirrors were re-opened at 16:04 UTC with standard powers. 

Images attached to this report
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×